# The distribution of interior branch lengths in phylogenetic trees under constant-rate birth-death models

### v1

@@ V Branch length distribution: interior branches
As a last step, we derive the distribution of internal branch lengths, i.e. of all branches that are not pendant. To this end, we calculate the joint probability of an interior branch starting at time $\tau_{s}$ and ending at $\tau_{e}<\tau_{s}$. As before, the probability per unit time of a birth event occurring at time $\tau_{s}$ is proportional to $\lambda\operatorname{\mathrm{e}}^{-(1-q)\tau_{s}}$. Focusing for now on only one of the two clades that started at $\tau_{s}$ (called $A$), the probability of it running until $\tau_{e}$ with only one extant descendant at the time of observation is $p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})$. The probability per small time interval of the branch ending at $\tau_{e}$ is equal to the probability per small time interval of a birth event at that time where none of the two emerging clades dies out. This is given by $\lambda(1-p_{0}(\tau_{e}))^{2}$.
Two cases have to be considered for the other clade (called $B$) emerging at $\tau_{s}$. With probability $p_{1}(\tau_{s})$ this clade becomes a pendant branch. In this case, a factor of two arises due to symmetry (either clade can be called $A$). An alternative scenario is the clade having more than one extant descendant. This happens with probability $p_{>1}(\tau_{s})=1-p_{0}(\tau_{s})-p_{1}(\tau_{s})$. In this case, the clade also has an interior branch starting at $\tau_{s}$ and therefore contributes to the branch length distribution. A factor of two accounts for this.
Putting these steps together, the relative probability of an interior branch running from $\tau_{s}$ to $\tau_{e}$ (up to a normalising factor found to be one) is given by
@@
Figure 2E compares this result to the empirical distribution of a tree reconstructed from a population of $7500$ individuals grown from a single individual with birth rate $\lambda=1$ and death rate $\mu=0.75$ as described above.
@@ VI Number of generations along branches
$\langle i\rangle(\tau_{s},\tau_{e})=2\lambda\int_{\tau_{e}}^{\tau_{s}}\operatorname{\mathrm{d}\!}\tau\,p_{0}(\tau)=2\left[\lambda\tau-\log\left(\lambda-\mu e^{-(\lambda-\mu)\tau}\right)\right]_{\tau_{e}}^{\tau_{s}}$ (15)
based on birth events occurring along a branch at rate $\lambda$ and the probability $p_{0}(\tau)$ that one given additional lineage has died out by time $\tau$ given by (2). The factor $2$ arises because either of the two offspring can be the one to die out. The number of generations is Poisson distributed (and hence fully characterised by the mean (15)), because successive birth events are statistically independent Bokma et al. (2012).
Together with the densities of branch lengths from the previous sections, we can now compute the probability distribution for the number of generations along a randomly chosen branch. In the case of pendant branches, it is given by
@@ VII Finite sampling
In many practical applications, the phylogenetic tree is not based on the entire extant population at the time of observation but on a subset of individuals that were sampled at that time. Reconstructing the phylogenetic tree from the sampled individuals changes the statistics of branch lengths derived above. A standard model for the process of sampling from the extant population is to choose extant individuals at the time of observation independently and with equal probability $\rho$ Yang and Rannala (1997); Stadler (2010); Stadler and Steel (2012). The preceding results on branch length distributions can be easily generalised to a sampling probability $\rho<1$ using the probabilities that a clade of a particular age has no (or one) member which is sampled at a particular time.
@@
To derive the probability density function of branch lengths under sampling, these probabilities can be used directly in the densities of pendant and interior branch lengths at full sampling ($\rho=1$), (10)-(13), which depend only on $p_{0}$, $p_{1}$, and the relative population sizes at different times. In particular, we get
@@
Note that a normalising factor of $\rho^{-1}$ applies to both densities of interior branches, while the normalising factor in the case of pendant branches again turns out to be one. The case of pendant branches extends the result of Stadler and Steel Stadler and Steel (2012), which was restricted to a regime of high death rate or high sampling probability $1\geq\mu/\lambda\geq 1-\rho$. The density function (22) gives the same result but also holds outside these limits.
Stadler Stadler (2009) found that the probability distribution of bifurcation times is not unique to every set of parameters, but that a birth-death process with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho$ yields the same distribution as a process with parameters $\rho^{\prime}$, $\lambda^{\prime}=\rho\lambda/\rho^{\prime}$, $\mu^{\prime}=\mu-\lambda(1-\rho/\rho^{\prime})$. We find the same for the distributions of branch lengths (21)-(23), as $\lambda^{\prime}-\mu^{\prime}=\lambda-\mu$, $\rho^{\prime}\lambda^{\prime}=\rho\lambda$ and $\lambda^{\prime}(1-\rho^{\prime})-\mu^{\prime}=\lambda(1-\rho)-\mu$ (which is the same argument as in Stadler (2009, 2013a)). The inference of both growth parameters and the sampling probability from the distribution of branch lengths is therefore not possible. However, the maximal sampling probability can be estimated: Given a branch distribution generated with (unknown) parameters $(\lambda,\mu,\rho)$, the parameter space of all sets of $(\lambda^{\prime},\mu^{\prime},\rho^{\prime})$ which lead to the same distribution is restricted to $\mu^{\prime}\geq 0$ and $0<\rho\leq 1$. With the transformation rules above, it is easy to see that these inequalities are only fulfilled if $0<\rho^{\prime}\leq\min(\rho\lambda/(\lambda-\mu),1)$. This means that for branch distributions generated with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho<1-\mu/\lambda$, there exists a maximal sampling probability $\rho^{\prime}<1$ for the family of alternative parameter sets.
The lengths of pendant and interior branches behave very differently under sampling. Both cases are depicted in Figure 3A. The rows correspond to different sampling probabilities $\rho$. Histograms show the branch length distributions of 100 simulations of trees with 300 sampled individuals, while the orange dashed lines depict the analytic results. Figure 3A shows in the second column the length distribution of pendant branches, which get longer with decreasing sampling probability $\rho$. On the other hand, interior branches quickly reach an asymptotic distribution as $\rho$ is decreased and then do not get any longer as $\rho$ is decreased further, see the first column of Figure 3A. The red dotted line depicts the asymptotic state, which can be derived as
@@
Figure 3: Pendant and interior branches at different sampling probabilities $\rho$. (A) We compare the analytical results for the density of the length of interior and pendant branches under sampling, (23) and (21) respectively, to numerical simulations. For the latter, we show the distribution of branches from 100 trees, each with 300 pendant branches. For every tree, we simulated the growth of a population with $\lambda=1$ and $\mu=0.75$ up to a size of $300/\rho$ and reconstructed the tree based on 300 randomly chosen individuals. (Fixing both population and sample size in this way is distinct from independent sampling of each individual. However, the difference vanishes with increasing population size.) Trees are grown in this way for different values of the sampling probability $\rho$, depicted in the different rows. The first column shows the distribution of interior branches, quickly reaching the asymptotic distribution (24) (red dotted line) as $\rho$ becomes small. The second column shows the distributions of pendant branch lengths, which shift to longer branches as $\rho$ decreases. (B) The distribution of the number of generations along interior branches in the asymptotic case of small sampling probabilities (25). Different colours represent different relative death rates $\mu/\lambda$. (C) A toy example of how sampling (here at sampling probability $\rho=1/2$) affects the length of pendant branches, see text.
Following the steps in Section VI using equations (21)-(23), we can compute the distribution of generations along branches of a sampled tree. Crucially, since the mean number of generations (equation (15)) along a branch of length $\tau$ converges to $2\mu\tau$ as the branch moves further away from the time of observation (as interior branches do when taking the limit $\rho\rightarrow 0$), we can compute the probability of finding $i$ generations along an interior branch in this limit
@@
This result is depicted for different relative death rates $\mu/\lambda$ in Figure 3B.
VIII Discussion
We have calculated the statistics of branch lengths (in calendar time) of phylogenetic trees generated by a constant-rate birth-death process and reconstructed from the population at a single moment in time. Our approach simplifies the derivation of the distribution of pendant branch lengths by Stadler and Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and allows to treat both branches at a given distance from the tree leaves as well as the branch length distribution of randomly chosen interior branches. We further generalised the distributions to the case of finite sampling. While we derived our distributions conditioned on the population size at the time of sampling, they turn out to be independent of that size. Other cases, such as conditioning on the first bifurcation time or on both bifurcation time and population size are still open for further investigation.
We found that pendant branches and interior branches behave very differently under sampling: only pendant branch lengths increase indefinitely as the sampling probability is decreased to zero, whereas the distribution of interior branches quickly reaches an asymptotic. This behaviour allows to place bounds on the model parameters (birth and death rates, sampling probability) on the basis of empirical trees.
@@ VIII Discussion
The behaviour of interior branch lengths also has consequences for the number of generations along the branches: even at low sampling probabilities, the number of generations along interior branches can be small — depending on the ratio of death and birth rates, but asymptotically independently of the sampling. Under cellular reproduction, when mutations occur at birth events, this means that the number of mutations along a branch can be shaped by the (few) birth events along that branch rather than accumulating continuously over time. The consequences and applications of this result will be explored in a future publication.

### v2

@@ V Branch length distribution: interior branches
As a next step, we derive the distribution of internal branch lengths, i.e., all branches that are not pendant. (We exclude the branch between the start of the birth-death process to the first bifurcation.) To this end, we calculate the joint probability of an interior branch starting at time $\tau_{s}$ and ending at $\tau_{e}<\tau_{s}$. As before, the probability per unit time of a birth event occurring at time $\tau_{s}$ is proportional to $\lambda\operatorname{\mathrm{e}}^{-(1-q)\tau_{s}}$. Focusing for now on only one of the two clades that started at $\tau_{s}$ (called $A$), the probability of it running until $\tau_{e}$ with only one extant descendant at the time of observation is $p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})$. The probability per small time interval of the branch ending at $\tau_{e}$ is equal to the probability per small time interval of a birth event at that time where none of the two emerging clades dies out. This is given by $\lambda(1-p_{0}(\tau_{e}))^{2}$.
Two cases must be considered for the other clade (called $B$) emerging at $\tau_{s}$. With probability $p_{1}(\tau_{s})$ this clade becomes a pendant branch. In this case, a factor of two arises due to symmetry (either clade can be called $A$). An alternative scenario is the clade having more than one extant descendant. This happens with probability $p_{>1}(\tau_{s})=1-p_{0}(\tau_{s})-p_{1}(\tau_{s})$. In this case, the clade also has an interior branch starting at $\tau_{s}$ contributing to the branch length distribution. A factor of two accounts for this.
Putting these steps together, the relative probability of an interior branch running from $\tau_{s}$ to $\tau_{e}$ (up to a normalization factor found to be one) is given by
@@
Figure 2E compares this result to the empirical distribution of a tree reconstructed from a population of age $T=25$ grown from a single individual with birth rate $\lambda=1$ and death rate $\mu=0.75$ as described above.
So far, we have again only considered the asymptotic case of $T\rightarrow\infty$. For finite $T$, the probability of $\tau_{s}>T$ is 00, resulting in a normalization factor of $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(a(1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))^{2}$ for (13).
@@ VI Number of generations along branches
$\langle i\rangle(\tau_{s},\tau_{e})=2\lambda\int_{\tau_{e}}^{\tau_{s}}\operatorname{\mathrm{d}\!}\tau\,p_{0}(\tau)=2\left[\mu\tau-\log\left(\lambda-\mu e^{-(\lambda-\mu)\tau}\right)\right]_{\tau_{e}}^{\tau_{s}}$ (15)
based on birth events occurring along a branch at rate $\lambda$ and the probability $p_{0}(\tau)$ that one given additional lineage has died out by time $\tau$ given by (2). The factor $2$ arises because either of the two offspring can be the one to die out. The number of generations is Poisson distributed (and hence fully characterized by the mean (15)), because successive birth events are statistically independent Bokma et al. (2012).
Together with the densities of branch lengths from the previous sections, we can now compute the probability distribution for the number of generations along a randomly chosen branch. In the case of pendant branches and asymptotically old trees, it is given by
@@ VII Finite sampling
Usually only a subset of individuals sampled from the population is available to construct the phylogenetic tree from and this sampling affects the statistics of branch lengths. To model the process of sampling from the extant population we follow the standard approach and choose extant individuals at the time of observation independently and with equal probability $\rho$ Yang and Rannala (1997); Stadler (2010); Stadler and Steel (2012). The preceding results on branch length distributions can be easily generalized to a sampling probability $\rho<1$ using the probabilities that a clade of a particular age has no (or one) member which is sampled at a particular time.
@@
Note that in this case, $p_{1}^{(\rho)}(\tau)=\rho p_{1}^{(\rho)(\tau_{e}=0)}(\tau)$, because we need to include the probability that the lineage running until $\tau_{e}=0$ is actually sampled.
To derive the probability density function of branch lengths under sampling, these probabilities can be used directly in the densities of pendant and interior branch lengths at full sampling ($\rho=1$), (10)-(13), which depend only on $p_{0}$, $p_{1}$, and the relative population sizes at different times. In particular, we obtain
@@
A normalizing factor of $\rho^{-1}$ applies to all three densities and is included in the expressions above. The result for pendant branches coincides with the result of Stadler and Steel Stadler and Steel (2012) at fixed number of samples. However, their result was restricted to a regime of high death rate or high sampling probability $1\geq\mu/\lambda\geq 1-\rho$. The density function (21) also holds outside these limits.
Stadler Stadler (2009) found that the probability distribution of bifurcation times is not unique to a particular set of parameters, but that a birth-death process with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho$ yields the same distribution as a process with parameters $\rho^{\prime}$, $\lambda^{\prime}=\rho\lambda/\rho^{\prime}$, $\mu^{\prime}=\mu-\lambda(1-\rho/\rho^{\prime})$. We find the same for the distributions of branch lengths (21)-(23), as $\lambda^{\prime}-\mu^{\prime}=\lambda-\mu$, $\rho^{\prime}\lambda^{\prime}=\rho\lambda$ and $\lambda^{\prime}(1-\rho^{\prime})-\mu^{\prime}=\lambda(1-\rho)-\mu$ (which is the same argument used in Stadler (2009, 2013a)). Therefore, inference of both growth parameters and the sampling probability from the distribution of branch lengths is impossible. However, an upper bound on the sampling probability can be determined: Given a branch distribution generated with (unknown) parameters $(\lambda,\mu,\rho)$, the parameter space of all sets of $(\lambda^{\prime},\mu^{\prime},\rho^{\prime})$ which lead to the same distribution is restricted to $\mu^{\prime}\geq 0$ and $0<\rho\leq 1$. With the transformation rules above, it is easy to see that these inequalities are only fulfilled if $0<\rho^{\prime}\leq\min(\rho\lambda/(\lambda-\mu),1)$. This means that for branch distributions generated with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho<1-\mu/\lambda$, there exists a maximal sampling probability $\rho^{\prime}<1$ for the family of possible parameter sets.
The lengths of pendant and interior branches behave very differently under sampling. Pendant and interior branch lengths are compared in Figure 3A. The rows correspond to different sampling probabilities $\rho$, the first column shows pendant, the second column interior branches. Histograms show the branch length distributions of $100$ simulations of trees, while the blue lines depict the analytic results. Pendant branches in the first column get longer with decreasing sampling probability $\rho$. On the other hand, interior branches quickly reach an asymptotic distribution as $\rho$ is decreased and then do not get any longer as $\rho$ is decreased further, see the second column of Figure 3A. The red dashed line depicts the asymptotic state, which can be derived as
@@
Figure 3: Pendant and interior branches at different sampling probabilities $\rho$. (A) We compare the analytical results for the probability density of the length of pendant and interior branches under sampling ((21) and (23) respectively, solid blue lines) to numerical simulations. For the latter, we show the distribution of branches from 100 trees. For every tree, we simulated the growth of a population with $\lambda=1$ and $\mu=0.75$ over time $T=40$, restarting when the population died out or the final population size was not within $2\times 10^{4}$ and $3\times 10^{5}$ (grey bars). The first row of plots shows the branch length distribution of the full trees. For the next row of plots, individuals were sampled with probability $\rho=0.1$ and the corresponding trees were reconstructed. For the third row, each of the previous samples was sampled again with probability $0.1$, corresponding to an overall sampling probability of $\rho=0.01$. The same was repeated for the forth row of plots, yielding $\rho=0.001$. Every tree contributes equally to the empirical distributions depicted. The first column shows the distributions of pendant branch lengths, which shift to longer branches as $\rho$ decreases. The second column of plots shows the distribution of interior branches, quickly reaching the asymptotic distribution (24) (red dashed line) as $\rho$ becomes small. (B) The distribution of the number of generations along interior branches in the asymptotic case of small sampling probabilities (25). Different colours represent different relative death rates $\mu/\lambda$. (C) A toy example of how sampling affects the length of pendant branches, see text.
Following Section VI and equations (21)-(23), we also compute the distribution of generations along branches of a sampled tree. Crucially, since the mean number of generations (equation (15)) along a branch of length $\tau$ converges to $2\mu\tau$ as the branch moves further away from the time of observation (as interior branches do when taking the limit $\rho\rightarrow 0$), we obtain the probability of finding $i$ generations along an interior branch in this limit
@@
This asymptotic result is depicted for different relative death rates $\mu/\lambda$ in Figure 3B.
VIII Branch length distributions conditioned on the population size at the time of observation
We now turn to the second statistical scenario, where the birth-death process is conditioned on the number $N$ of extant individuals rather than a particular time of observation (scenario ii) above). This case is harder to treat than scenario i), because with a constraint on the final number of individuals, the dynamics of different clades are in general correlated. We set up a framework to calculate branch length distributions based on tracking the dynamics of the birth-death process first backwards in time from the time of observation to the emergence of a particular branch and then again forward in time to the time of observation. For pendant branches, this approach recovers the results of Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012), but it also allows treating interior branches.
We first illustrate the method with a toy problem: We calculate the distribution of the length of pendant branches when the number of individuals at the time of observation is $N=2$. A key tool is the probability that a birth-death process with birth rate $\lambda$ and death rate $\mu$, starting with $n$ individuals at a given time has $N$ individuals a time interval $t$ later, denoted $p_{N,n}^{\lambda,\mu}(t)$.
Figure 4: Pendant branch formation for populations of size $N=2$: The red branch marks the pendant branch arising from a birth event at time $\tau$ in the past. All other $n-1$ individuals that were alive at time $\tau$ need to die out before the population is observed, while the second descendant from the birth event needs to survive with clade size one. The arrows on the right illustrate the conditional probabilities of having $N$ extant descendants given that $n$ individuals at time $\tau$ before observation (forwards process) and the probability that given $N$ observed individuals there were $n$ individuals at time $\tau$ before observation (backwards process).
First we go backwards in time from the time of observation to the time $\tau$. Conditioned on the number of individuals at the time of observation being $N$, the statistics of the number of individuals before the time of observation is described by a birth-death process with time running backwards and with the rates of birth and death exchanged, so the rate of birth is $\mu$ and the rate of death is $\lambda$ Ignatieva et al. (2020). The probability of having $n$ individuals at time $\tau$ before the time of observation is thus $p_{n,N}^{\mu,\lambda}(\tau)$. The number of individuals $n$ will eventually be summed over.
With $n$ individuals at time $\tau$, the expected number of birth events at time $\tau$ (per small time interval) leading to pendant branches, conditioned on $N=2$ individuals at the time of observation is
$\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)/p_{2,n}^{\lambda,\mu}(\tau)\ .$ (26)
$\lambda n$ gives the probability per small time interval that one of the $n$ individuals divides during this small interval, $p_{1}^{2}(\tau)$ is the probability that the two resulting offspring each have one extant offspring at the time of observation, leading to pendant branches of length $\tau$. $p_{0}^{n-1}(\tau)$ is the probability that all other $n-1$ individuals die out by the time of observation, leading to a population of size $N=2$. The denominator arises from the conditioning on $N=2$ individuals. What we have done is first asking about the statistics of the number of individuals at a certain time $\tau$ in the past, given $N$ individuals at some time of observation, and then asking about the probability that a tree with a particular feature (here a pendant branch that emerged at time $\tau$) arises from those forward-time trajectories ending in $N$ individuals at the time of observation. For an illustration, see Figure 4.
Up to a normalizing constant, the probability density function of the pendant branch lengths is thus
$\sum_{n=1}^{\infty}\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)\frac{p_{n,2}^{\mu,\lambda}(\tau)}{p_{2,n}^{\lambda,\mu}(\tau)}\ .$ (27)
In Appendix .1 we show how this expression evaluates to $2\lambda\frac{p_{1}^{2}}{1-p_{0}}=2\lambda\frac{(\lambda-\mu)^{3}\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{3}}$ and a normalisation constant of one. We find the same result for a final population size $N\geq 2$. This reproduces the result by Stadler and collaborators on the distribution of pendant branches conditioned on the size $N$ of the population at the time of observation. Curiously, this distribution also coincides with the distribution of pendant branch lengths under scenario i) in the limit of large $\tau$ (equation (10)).
The same approach can also be used to calculate the distribution of interior branch lengths conditioned on the population size. Again we use the birth-death process with time running backwards to compute the probability that at a time $\tau_{s}$ in the past there were $n$ individuals, one of which divided establishing a new branch. Running time forwards again and conditioning on the final population size $N$, we calculate the probability that the new branch ends at $\tau_{e}$ in another birth event, with both descendants having extant offspring. The conditions that lead to a branch of a particular length are the same manner as in scenario i), the difference is the conditioning on the final population size in two directions of time. The calculation is given in Appendix .2.
The resulting distribution of interior branch lengths under scenario ii) depends on $N$. However, the dependency quickly vanishes with $\left(\frac{\lambda}{\mu}p_{0}(\tau_{s})\right)^{N-1}$ when increasing the population size. The asymptotic distribution for large $N$ coincides with the interior branch length distribution under scenario i) in the limit of large times, see Appendix .2.
IX Discussion
We have calculated the statistics of branch lengths (in calendar time) of phylogenetic trees generated by a constant-rate birth-death process and reconstructed from the population at a single moment in time, with a particular focus on how branch lengths change when the population is sampled at different sampling probabilities. We look at two distinct statistical scenarios: sampling at a particular time since the birth-death process started and sampling when the population has a given size $N$.
@@ IX Discussion
For the first scenario, we generalize the results of Paradis Paradis (2016) to a finite sampling probability. We find that pendant branches and interior branches behave very differently under sampling: only pendant branch lengths increase indefinitely as the sampling probability is decreased to zero, whereas the distribution of interior branches quickly reaches an asymptotic. This behaviour allows placing bounds on the model parameters (birth and death rates, sampling probability) on the basis of empirical trees. A second consequence of this behaviour under sampling concerns the number of generations along interior branches: even at low sampling probabilities, the number of generations along interior branches can be small — depending on the ratio of death and birth rates, but asymptotically independently of the sampling. Under cellular reproduction, when mutations occur at birth events, this means that the number of mutations along a branch can be shaped by the (few) birth events along that branch rather than accumulating continuously over time. The consequences and applications of this result will be explored in a future publication.
For the second scenario, a fixed given final population size $N$, we develop a statistical framework based on first going backwards in time from a fixed population size, and then forwards again while conditioning on that population size. The latter step entails calculating the probability of a population with a given final size and a particular feature of its phylogenetic tree (like a branch of a given length), relative to the probability of a population with a certain size. The combination of going backwards and forwards in time is similar in spirit to the celebrated Keldysh––Schwinger formalism in non-equilibrium quantum physics Kamenev (2011). Our approach simplifies the derivation of the distribution of pendant branch lengths by Stadler and Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and allows also to treat the length distribution of interior branches.
The interior branch distributions of the asymptotic cases of both scenarios (large times and large population sizes) turn out to coincide. In case of pendant branches, the asymptotic distribution in the limit of large times since the start of the birth-death process coincides with the distribution in trees conditioned on population size N, even when N is not large, as the latter distribution is independent of $N$ Stadler and Steel (2012).